# Authorship Verification based on the Likelihood Ratio of Grammar Models

### v1

@@ 8       $(t\mbox{}_{1},t\mbox{}_{2},\ldots,t\mbox{}_{m})\leftarrow S\mbox{}$ for $j\leftarrow 1$ to $m$ do
In addition to this property, the $\lambda_{G}$ value is also effectively an uncalibrated log-likelihood ratio score that includes information about both similarity and typicality and that can then be calibrated into a forensic log-likelihood ratio using a calibration method such as a logistic regression. We denote a calibrated $\lambda_{G}$ log-likelihood ratio that expresses the strength of evidence in a forensic context with the upper case lambda ($\Lambda_{G}$). To perform this calibration we use a training corpus containing Y- and N-cases to estimate a logistic regression calibration model [110, 49] that turns each $\lambda_{G}$ into $\Lambda_{G}$. Here, $\Lambda_{G}$ quantifies the strength of the linguistic evidence for each of the two alternative forensic propositions. More details on the Likelihood Ratio Framework for forensic science and its evaluation is described in Section 5.4. Based on $\Lambda_{G}$, we classify a verification case as Y if $\Lambda_{G}>0$, else as N.
Finally, we note that another important aspect that is not discussed in detail in this section is the fact that several of the key components of LambdaG are consistent with Cognitive Linguistics theories. More details about this aspect of LambdaG are left to Section 4.
@@ 3 Results
Algorithm 1 was applied to a set of twelve corpora that simulate various real-life AV problems and scenarios, described in Section 5. The full results are presented in Table 1, where the performance of each method is evaluated using five standard performance measures for classification: Accuracy, the Area Under the Curve (AUC), F1, Precision, and Recall. These are values that were calculated using a threshold $\theta$, which is derived from $\Lambda_{G}$. The split between training and test datasets for each corpus tested is shown in Table 4. To assess the performance of $\Lambda_{G}$, instead, the $C_{\text{llr}}\mbox{}$ measure is used, which expresses the amount of information in the likelihood ratio. This metric is explained in Section 5.4.
In Table 1, LambdaG is compared against four well-known and tested AV methods: IM, COAV, TAVeer and a neural language model using a Siamese network, SiamBERT, which is explained in Section 5.5. IM and COAV were run on POSNoise pre-processed corpora and they are therefore content-agnostic, similarly to LambdaG. TAVeer was run on the unprocessed corpora but because the method only measures topic-agnostic features, this is also comparable to LambdaG. Since SiamBERT uses a pre-trained model on data that it is not POSNoise-processed, fine-tuning it on POSNoise-processed data would be an unfair comparison. Equally, however, SiamBERT fine-tuned on unprocessed data risks using content information and is also, in a different way, an unfair comparison. For this reason, we present the results of both versions. The results of SiamBERT on POSNoise-processed data is indicated with ${}^{\,(\bm{\dagger})}$.
The results in Table 1 show that LambdaG is superior to the other AV topic-agnostic baselines in all cases. LambdaG is also the best performing method when compared to SiamBERT despite the fact that any correlation between author and topic could have been exploited in each corpora. The only exception is for $\mathcal{C}_{\mathrm{AllNews}}$, where instead LambdaG comes second to SiamBERT. We do not have an explanation for this result but we advance the plausible hypothesis that $\mathcal{C}_{\mathrm{AllNews}}$ contains data with a significant author/topic bias such that the task is made easier when content words are included. This hypothesis is supported by the extremely poor performance of ${}^{\,(\bm{\dagger})}$ on $\mathcal{C}_{\mathrm{Stack}}$, which is a corpus explicitly designed to penalize AV methods that exploit topic information.
@@
Table 1: Evaluation results of LambdaG and the four baseline methods. ${}^{\,(\bm{\dagger})}$ indicates that the method was applied to the original (i. e., non-POSNoise-processed) corpora. In terms of accuracy and AUC, bold and underlined values represent the best and second-best results, respectively.
@@
Table 2 instead shows the values of $C_{\text{llr}}$ and $C_{\text{llr}}^{\text{min}}$ for both $\lambda_{G}$ and $\Lambda_{G}$, since both are log-likelihood ratios. A significant discrepancy is observed between the $C_{\text{llr}}$ and the $C_{\text{llr}}^{\text{min}}$ values for $\lambda_{G}$, indicating that $\lambda_{G}$ is uncalibrated. This means that although higher values of $\lambda_{G}$ do correctly correspond to Y-cases, the scale of variation does not reflect the expectations of a perfectly calibrated system, where $\lambda_{G}=0$ means an inconclusive result, a positive value suggests a Y-case, and a negative value suggests an N-case. When $\lambda_{G}$ is turned into $\Lambda_{G}$ by fitting a logistic regression on training data, however, the lower values of $C_{\text{llr}}\mbox{}(\Lambda_{G})$ are evidence that $\Lambda_{G}$ is well calibrated.
All results in Table 1 and Table 2 were obtained using LambdaG with parameters $r=100$ and $N=10$. Figure 1 shows how changes in these hyper-parameters affects Accuracy, revealing that the method is robust in the choice of these hyper-parameters. It is clear from the graph that $r=1$, meaning that only one reference Grammar Model $G_{\text{ref}}$ is extracted from the reference corpus, is not enough to achieve optimal results. However, in most corpora $r=100$ is not needed as performance stabilize at $r=30$.
Regarding the other hyperparameter $N$, the order of the $n$-gram model, the value $N=10$ was chosen to ensure that grammatical dependencies are captured even for long sentences and Figure 1 does suggest that this assumption is correct. For example, for a corpus such as $\mathcal{C}_{\mathrm{PJ}}$ consisting of chat messages or $\mathcal{C}_{\mathrm{Enron}}$ consisting of emails, $N=2$ is enough. In contrast, for the academic papers in $\mathcal{C}_{\mathrm{ACL}}$ and the news in $\mathcal{C}_{\mathrm{AllNews}}$ a larger order of the model clearly has an advantage. In all corpora, however, there is no gain in increasing $N$ to 20.
Figure 1: Accuracy results for different numbers of repetitions $r$ and orders of the $n$-gram model $N$ for each corpus.
Figure 2: The loss in Accuracy (top) and $C_{\text{llr}}$ (bottom) results for cross-corpus comparison, i.e. evaluating on Base Corpus while using reference texts $\mathbb{D}_{\text{ref}}$ from Reference Corpus. Diagonal bold values denote the original Accuracy and $C_{\text{llr}}$, respectively. Darker shades denote a greater loss in performance.
Another important robustness test that we performed is the selection of the reference data. Reliance on external data is a potential limitation for any binary extrinsic AV methods, like IM. As explained in Section 1, IM, suffers a substantial loss in performance if the impostors are selected from another genre. LambdaG also has the crucial requirement that a reference corpus $\mathbb{D}_{\text{ref}}$ is collected and it is presumably important that $\mathbb{D}_{\text{ref}}$ is as compatible as possible, especially in genre but generally speaking in as many parameters as possible, with $\mathbb{D}_{\mathcal{A}}$ and $\mathbb{D}_{\mathcal{U}}$. To test the robustness of LambdaG to variations in $\mathbb{D}_{\text{ref}}$ type we ran 144 tests by pairing each of the twelve corpora to every other corpora. For example, we performed the analysis for $\mathcal{C}_{\mathrm{Enron}}$ but we used as $\mathbb{D}_{\text{ref}}$ $\mathcal{C}_{\mathrm{ACL}}$. We then calculated Accuracy and $C_{\text{llr}}$ for all combinations and we produced a plot for each metric displaying the resulting drop in performances when using a different $\mathbb{D}_{\text{ref}}$ (Figure 2).
@@
The diagonal of Figure 2 (top) shows the same results reported in Table 1, which function as the benchmark of performance for each row of the plot, corresponding to the corpus being tested. Overall the results suggest that LambdaG is extremely robust in many cases of genre variation, including some unrealistic genre combinations. This result is unexpected given how IM is instead very much dependent on the set of impostors [41]. For example, cross-corpus comparisons such as using $\mathcal{C}_{\mathrm{Yelp}}$ as reference for problems in $\mathcal{C}_{\mathrm{Amazon}}$ lead to very small losses in Accuracy and this is clearly explainable by the fact that both corpora contain reviews. However, certain results are more surprising, such as that using the academic papers from $\mathcal{C}_{\mathrm{ACL}}$ as reference data for problems in a corpus of chat logs such as $\mathcal{C}_{\mathrm{PJ}}$ leads to a performance loss of only 0.27 Accuracy, which is substantial but much less than expected given the disparity. This is a comparison that a human analyst would be unlikely to perform and therefore it is a real test of the lower limits in terms of robustness to reference corpora.
@@
Looking at Figure 2 (top) row-wise, the most notable pattern is that the two corpora that suffer the most performance loss overall are $\mathcal{C}_{\mathrm{PJ}}$ and $\mathcal{C}_{\mathrm{Stack}}$. For $\mathcal{C}_{\mathrm{PJ}}$ the explanation is likely to be that it is the only corpus with sentences mostly corresponding to chat messages, and therefore any match of this corpus with other corpora results in at least this incompatibility. The chat logs in $\mathcal{C}_{\mathrm{PJ}}$ are also the texts with the lowest level of formality and highest degree of linguistic freedom among all corpora. For $\mathcal{C}_{\mathrm{Stack}}$ this result could be due to its hybrid genre character (forum posts on academic topics) or it could simply be due the use of several tags to mask mathematical expressions or computer code. When looking at Figure 2 column-wise, instead, the most notable pattern is that using $\mathcal{C}_{\mathrm{ACL}}$ or $\mathcal{C}_{\mathrm{AllNews}}$ as reference corpora tends to affect performance more drastically for the majority of corpora, which again could be explained by their relatively higher degree of formality compared to the other corpora. All in all this analysis suggests that, for $\mathcal{C}_{\mathrm{PJ}}$ and $\mathcal{C}_{\mathrm{Stack}}$ the reference corpus has a moderate impact on performance while for all other corpora it is only the use of $\mathcal{C}_{\mathrm{ACL}}$ or $\mathcal{C}_{\mathrm{AllNews}}$ as reference where this modest impact can be observed. For all other cross-corpus comparisons for most corpora the effect of the reference corpus is tolerable.
@@
We also present the results for the $C_{\text{llr}}$ metric in Figure 2 (bottom). As explained in Section 5.4, a $C_{\text{llr}}$ value close to or equal to one means that the system is not providing any useful information for the case, while a $C_{\text{llr}}$ value higher than 1 means that results tend to be misleading (e.g. negative $\Lambda_{G}$ for a Y-problem). None of the cross-genre comparisons produce $C_{\text{llr}}$ higher than one, meaning that in all cases the values of $\Lambda_{G}$ are not misleading the analyst but can be, at worst, inconclusive with $\Lambda_{G}\approx 0$. These results support the principle that an analyst should always compile a reference data set that is as close as possible to the case data [52]. However, the results do also suggest that LambdaG is stable to small variations in genre.
@@
The final test we carried out is a comparison of LambdaG as implemented in Algorithm 1 vs. LambdaG without applying the POSNoise function. The rationale for this test is to understand whether more authorial information can be extracted and exploited using this method beyond just grammatical structure. The results in Table LABEL:tab:EvalResultsPosnoise show that, regarding Accuracy and AUC, in all cases the proposed version of Algorithm 1 with the POSNoise pre-processing is better performing than a version of the algorithm without this pre-processing. The difference in Accuracy is not large but the method is clearly superior when the focus is entirely on grammar to the extent that adding information about meaning instead leads to adding noise in the authorial signal.
As explained in Section 5.3, the advantage of LambdaG is that it can be better interpreted than other methods. As a first step, the analyst can produce a ranking of all sentences in $\mathcal{D}_{\mathcal{U}}$ and then, additionally, each token can be color-coded according to its value of $\lambda_{G}$, similarly to COAV. Some of these visualizations explaining the features discovered are shown in Section 6.
@@
$\mathcal{C}_{\mathrm{Enron}}$
0.031 0.039 0.039 -0.017 0.083 4 -4 1 -1
$\mathcal{C}_{\mathrm{Wiki}}$
0.026 0.003 0.029 0.022 0.036 4 -4 -2 2
$\mathcal{C}_{\mathrm{Stack}}$
0.005 0.001 -0.002 0.047 -0.044 -5 5 -6 6
$\mathcal{C}_{\mathrm{ACL}}$
0.057 0.050 0.031 0.106 -0.050 -7 7 -23 23
$\mathcal{C}_{\mathrm{PJ}}$
0.009 0.005 0.010 0.001 0.019 3 -3 0 0
$\mathcal{C}_{\mathrm{Apric}}$
0.077 0.042 0.084 0.038 0.129 22 -22 -4 4
$\mathcal{C}_{\mathrm{Trip}}$
0.037 0.049 0.048 0.022 0.071 17 -17 -1 1
$\mathcal{C}_{\mathrm{Yelp}}$
0.057 0.046 0.063 0.043 0.083 20 -20 -7 7
$\mathcal{C}_{\mathrm{IMDB}}$
0.080 0.101 0.091 0.070 0.108 87 -87 -42 42
$\mathcal{C}_{\mathrm{Blog}}$
0.063 0.038 0.066 0.048 0.083 75 -75 -38 38
$\mathcal{C}_{\mathrm{Amazon}}$
0.046 0.030 0.046 0.041 0.050 60 -60 -49 49
$\mathcal{C}_{\mathrm{AllNews}}$
0.043 0.049 0.059 0.034 0.076 101 -101 -15 15

### v2

@@ 26       $(t\mbox{}_{1},t\mbox{}_{2},\ldots,t\mbox{}_{z})\leftarrow S\mbox{}_{i}$
In addition to this property, the $\lambda_{G}$ value is also effectively an uncalibrated log-likelihood ratio score that includes information about both similarity and typicality and that can then be calibrated into a forensic log-likelihood ratio using a calibration method such as a logistic regression. We denote a calibrated $\lambda_{G}$ log-likelihood ratio that expresses the strength of evidence in a forensic context with the upper case lambda ($\Lambda_{G}$). To perform this calibration we use a training corpus containing Y- and N-cases to estimate a logistic regression calibration model [130, 62] that turns each $\lambda_{G}$ into $\Lambda_{G}$. Here, $\Lambda_{G}$ quantifies the strength of the linguistic evidence for each of the two alternative forensic propositions. More details on the Likelihood Ratio Framework for forensic science and its evaluation is described in Section 6.4. Based on $\Lambda_{G}$, we classify a new verification case $c_{\textrm{new}}$ as Y if $\Lambda_{G}>0$ (that is, if $\text{sigmoid}(\Lambda_{G})>0.5$), else as N.
@@ 4 Results
Algorithm 1 was applied to a set of twelve corpora that simulate various real-life AV problems and scenarios, described in Section 6. The complete results are presented in Table LABEL:tab:EvalResults, where the performance of each method is evaluated using five standard performance measures for classification: Accuracy, the Area Under the Curve (AUC), $F_{\text{1}}$, Precision, and Recall. These are values that were calculated using a threshold $\theta$, which is derived from $\Lambda_{G}$. The split between training and test datasets for each corpus tested is shown in Table 5. To further assess the performance of $\Lambda_{G}$, the $C_{\text{llr}}\mbox{}$ measure is also used, which is a typical measure of accuracy for a likelihood ratio. This metric is explained in Section 6.4. When running the analysis on the training corpora, the reference corpus $\mathbb{D}_{\text{ref}}$ was the remaining of the corpus after taking out the authors of $\mathcal{D}_{\mathcal{A}}$ and $\mathcal{D}_{\mathcal{U}}$. When analyzing the test corpora, the reference corpus was simply the corresponding training corpus. A robustness test using different corpora as reference data was also carried out (see Figure 4). Given that LambdaG is non-deterministic, each analysis was repeated five times, and the results presented in Table LABEL:tab:EvalResults are the average of these five repetitions.
In Table LABEL:tab:EvalResults, LambdaG is compared with seven well-known and tested baseline AV methods: IM, COAV, TAVeer, ADHOMINEM, FeVecDiff (LR), FeVecDiff (MLP), and our own developed baseline SiamBERT, which is explained in Section 6.6. IM, COAV and FeVecDiff (LR and MLP) were run on POSNoise pre-processed corpora and are therefore content-agnostic, similarly to LambdaG. TAVeer, on the other hand, was applied to the unprocessed corpora. However, since this method is inherently designed to consider only topic-agnostic features [^7] Note that both TAVeer and POSNoise rely on the same list of topic-agnostic words and phrases shown in Table 6., its results remain comparable to those of LambdaG. With regard to SiamBERT and ADHOMINEM, we wish to point out that both methods are based on pre-trained neural network models, which were originally trained on POSNoise-unprocessed text data. Using pre-trained models is essential because training neural networks from scratch requires datasets that are larger than our AV corpora. Instead, we therefore fine-tuned the original pre-trained models on our AV corpora. Since pre-training was performed on POSNoise unprocessed text, the standard approach would be to fine-tune on similarly unprocessed AV corpora (denoted as ${}^{\,(\bm{\dagger})}$ and ${}^{\,(\bm{\dagger})}$ in Table LABEL:tab:EvalResults). However, this could introduce topic-related distortions, making the results less comparable with the other AV methods used. To mitigate this, we also fine-tuned and applied SiamBERT and ADHOMINEM on our POSNoise pre-processed corpora.
@@
Table 1: Evaluation results of LambdaG and the seven introduced baselines. $(\bm{\dagger})$ is marked next to a pre-trained neural network approach to indicate that it was fine-tuned using the original (i. e., non-POSNoise-processed) corpora. In terms of Accuracy and AUC, bold and underlined values represent the best and second-best results, respectively. For better readability, the confusion matrix outputs for each tested method have been grayed out.
@@
ADHOMINEM 0.506 0.511 0.572 0.505 0.660 879 452 862 469
${}^{\,(\bm{\dagger})}$ 0.551 0.549 0.610 0.539 0.702 935 396 799 532
In addition to Table LABEL:tab:EvalResults, Figure 2 shows 95% confidence intervals for Accuracy, calculated using the classical standard error estimate for a proportion, under the simplifying assumption of independent and identically distributed (i.i.d.) test data:
$\text{SE}=\sqrt{\frac{1}{n}\Bigl{(}\text{Accuracy}\times(1-\text{Accuracy})}\Bigr{)}$ (5)
Here, $n$ indicates the total number of tests performed. Even though there is no obvious reason to take the i.i.d. assumption seriously due to the non-trivial sampling process of our test datasets (see Section 6.1) and the relatively high number of potential comparisons involved, we can consider the resulting standard errors as a rough indication of the potential size of statistical fluctuations and an assessment of whether the observed performance improvements really suggest the presence of actual effects.
Figure 2: 95% Confidence Intervals for Accuracy.
When considering the overall pattern of results, we can conclude that LambdaG is superior to the other AV baselines as LambdaG is the only method that is consistently ranked at the top for Accuracy for most corpora. This consistency across many corpora representing different AV scenarios in terms of lengths and genres is a strong suggestion that this effect is genuine.
When judging the robustness of this difference by looking at the overlap of the confidence intervals, LambdaG’s superiority is clear for $\mathcal{C}_{\mathrm{Yelp}}$, $\mathcal{C}_{\mathrm{Blog}}$ and $\mathcal{C}_{\mathrm{Amazon}}$. IM is the only method that tends to be a close second in many corpora, often with substantial overlap in the error bars with LambdaG ($\mathcal{C}_{\mathrm{Enron}}$, $\mathcal{C}_{\mathrm{Wiki}}$, $\mathcal{C}_{\mathrm{ACL}}$, $\mathcal{C}_{\mathrm{PJ}}$, $\mathcal{C}_{\mathrm{Apric}}$, and $\mathcal{C}_{\mathrm{IMDB}}$). For $\mathcal{C}_{\mathrm{ACL}}$, although LambdaG is ranked first, there is a substantial overlap in the confidence intervals with both IM and COAV but all these methods are clearly superior to all the neural network approaches. There are only two corpora for which LambdaG is not ranked as first: $\mathcal{C}_{\mathrm{Stack}}$ and $\mathcal{C}_{\mathrm{AllNews}}$. For $\mathcal{C}_{\mathrm{Stack}}$, the best method is FeVecDiff (LR) but LambdaG is a close second, with the confidence interval substantially overlapping with FeVecDiff (MLP). For $\mathcal{C}_{\mathrm{AllNews}}$, LambdaG comes second to ${}^{\,(\bm{\dagger})}$ and this difference is significant. However, it is important to note that ${}^{\,(\bm{\dagger})}$ could have exploited any topic bias in the data. For all other corpora, LambdaG always performs better than ${}^{\,(\bm{\dagger})}$ and ${}^{\,(\bm{\dagger})}$, despite the fact that any correlation between author and topic could have been exploited in each corpora. We do not have an explanation for why ${}^{\,(\bm{\dagger})}$ outperforms all other methods only in $\mathcal{C}_{\mathrm{AllNews}}$, but we advance the hypothesis that $\mathcal{C}_{\mathrm{AllNews}}$ contains data with a significant author/topic bias such that the task is made easier when content words are included. This hypothesis is supported by the extremely poor performance of ${}^{\,(\bm{\dagger})}$ on $\mathcal{C}_{\mathrm{Stack}}$, which is a corpus explicitly designed to penalize AV methods that exploit topic information. Table 2 instead shows the values of $C_{\text{llr}}$ and $C_{\text{llr}}^{\text{min}}$ for both $\lambda_{G}$ and $\Lambda_{G}$, since both are log-likelihood ratios. A significant discrepancy is observed between the $C_{\text{llr}}$ and the $C_{\text{llr}}^{\text{min}}$ values for $\lambda_{G}$, indicating that $\lambda_{G}$ is uncalibrated. This means that although higher values of $\lambda_{G}$ do correctly correspond to Y-cases, the scale of variation does not reflect the expectations of a perfectly calibrated system, where $\lambda_{G}=0$ means an inconclusive result, a positive value suggests a Y-case, and a negative value suggests an N-case. When $\lambda_{G}$ is turned into $\Lambda_{G}$ by fitting a logistic regression on training data, however, the lower values of $C_{\text{llr}}\mbox{}(\Lambda_{G})$ are evidence that $\Lambda_{G}$ is well calibrated.
@@
All results shown in Table LABEL:tab:EvalResults and Table 2 were obtained with LambdaG using the hyperparameters $r=100$ and $N=10$. Figure 3 shows how changes in these hyperparameters affects Accuracy, revealing that the method is robust in the choice of these hyperparameters. It is clear from the graph that $r=1$, meaning that only one reference Grammar Model $G_{\text{ref}}$ is extracted from the reference corpus, is not enough to achieve optimal results. However, in most corpora $r=100$ is not needed as performance stabilize at $r=30$. Regarding the other hyperparameter $N$, the order of the $n$-gram model, the value $N=10$ was chosen to ensure that grammatical dependencies are captured even for long sentences and Figure 3 does suggest that this assumption is correct. For example, for a corpus such as $\mathcal{C}_{\mathrm{PJ}}$ consisting of chat messages or $\mathcal{C}_{\mathrm{Enron}}$ consisting of emails, $N=2$ is enough. In contrast, for the academic papers in $\mathcal{C}_{\mathrm{ACL}}$ and the news in $\mathcal{C}_{\mathrm{AllNews}}$ a larger order of the model clearly has an advantage. In all corpora, however, there is no gain in increasing $N$ to 20.
Figure 3: Variation in Accuracy depending on the number of repetitions, $r$, and orders of the $n$-gram model, $N$.
Another important robustness test that we performed is the selection of the reference data. Reliance on external data is a potential limitation for all binary extrinsic AV methods such as IM. As explained in Section 1, IM, suffers a substantial loss in performance if the impostors are selected from another genre. LambdaG has the same crucial requirement that a reference corpus $\mathbb{D}_{\text{ref}}$ is collected and it is presumably important that $\mathbb{D}_{\text{ref}}$ is as compatible as possible, especially in genre, with $\mathcal{D}_{\mathcal{A}}$ and $\mathcal{D}_{\mathcal{U}}$. To test the robustness of LambdaG to variations in $\mathbb{D}_{\text{ref}}$ type we ran 144 tests by pairing each of the twelve corpora to every other corpora. For example, we performed the analysis for $\mathcal{C}_{\mathrm{Enron}}$ but we used as $\mathbb{D}_{\text{ref}}$ $\mathcal{C}_{\mathrm{ACL}}$. We then calculated Accuracy and $C_{\text{llr}}$ for all combinations and we produced a plot for each metric displaying the resulting drop in performances when using a different $\mathbb{D}_{\text{ref}}$ (Figure 4). The diagonal of Figure 4 (top) shows the same results reported in Table LABEL:tab:EvalResults, which function as the benchmark of performance for each row of the plot, corresponding to the corpus being tested.
@@
These results show that, as expected, the performance of LambdaG can be strongly affected by genre variations in the reference corpus. However, the performance loss overall is not as dramatic as it would be expected for a method such as IM [51]. Firstly, for reasonable cross-corpus comparisons such as using $\mathcal{C}_{\mathrm{Yelp}}$ as reference for problems in $\mathcal{C}_{\mathrm{Amazon}}$, where both corpora are reviews, the loss in Accuracy is small. Secondly and more surprisingly, ‘unreasonable’ comparisons such using the academic papers from $\mathcal{C}_{\mathrm{ACL}}$ as reference data for problems in a corpus of chat logs such as $\mathcal{C}_{\mathrm{PJ}}$ leads to a reduction of, relatively speaking, only 0.27 Accuracy. Such a loss is clearly substantial but the expectation given the literature is that the method should perform at random in such a case. This is a comparison that a human analyst would be unlikely to perform and therefore it is a real test of the lower limits in terms of robustness to reference corpora.
@@
Examining Figure 4 (top) row-wise, the most notable pattern is that the two corpora that suffer the most performance loss overall are $\mathcal{C}_{\mathrm{PJ}}$ and $\mathcal{C}_{\mathrm{Stack}}$. For $\mathcal{C}_{\mathrm{PJ}}$ the explanation is likely to be that it is the only corpus with sentences corresponding to chat messages, and therefore any match of this corpus with other corpora suffers from this important difference. The chat logs in $\mathcal{C}_{\mathrm{PJ}}$ are also the texts with the lowest level of formality and highest degree of linguistic freedom among all corpora. For $\mathcal{C}_{\mathrm{Stack}}$ this result could be due to its hybrid genre character (forum posts on academic topics) or it could simply be due the use of several tags to mask mathematical expressions or computer code. When looking at Figure 4 column-wise, instead, the most notable pattern is that using $\mathcal{C}_{\mathrm{ACL}}$ or $\mathcal{C}_{\mathrm{AllNews}}$ as reference corpora tends to affect performance more drastically for the majority of corpora, which again could be explained by their relatively higher degree of formality compared to the other corpora. All in all, this analysis suggests that, for $\mathcal{C}_{\mathrm{PJ}}$ and $\mathcal{C}_{\mathrm{Stack}}$ the reference corpus has a moderate impact on performance, while for all other corpora it is only the use of $\mathcal{C}_{\mathrm{ACL}}$ or $\mathcal{C}_{\mathrm{AllNews}}$ as a reference where this modest impact can be observed. For most other cross-corpus comparisons, the effect of using a different reference corpus is tolerable.
@@
We also present the results for the $C_{\text{llr}}$ metric in Figure 4 (bottom). As explained in Section 6.4, a $C_{\text{llr}}$ value close to or equal to one means that the system is not providing any useful information, while a $C_{\text{llr}}$ value higher than 1 means that results tend to be misleading (e. g., negative $\Lambda_{G}$ for a Y-problem). None of the cross-genre comparisons produce $C_{\text{llr}}$ higher than one, meaning that in all cases, on average, the values of $\Lambda_{G}$ are not misleading the analyst but can be, at worst, inconclusive with $\Lambda_{G}\approx 0$. These results support the principle that an analyst should always compile a reference data set that is as close as possible to the case data [65]. However, the results do also suggest that LambdaG is certainly stable to small variations in genre or reference corpus composition.
Figure 4: The loss in Accuracy (top) and $C_{\text{llr}}$ (bottom) results for cross-corpus comparison, i. e., evaluating on Base Corpus while using reference texts $\mathbb{D}_{\text{ref}}$ from Reference Corpus. Diagonal bold values denote the original Accuracy and $C_{\text{llr}}$, respectively. Darker shades denote a greater loss in performance.
@@
The final tests we carried out relate to the contribution of POSNoise to the algorithm. We firstly performed a comparison of LambdaG as implemented in Algorithm 1 vs. LambdaG without applying the POSNoise function. The rationale for this test is to understand whether more authorial information can be extracted and exploited using this method beyond just grammatical structure. The results in Table 3 show that, regarding Accuracy and AUC, in all cases the proposed version of Algorithm 1 with the POSNoise pre-processing is better performing than a version of the algorithm without this pre-processing. The difference in Accuracy is not large but it is systematic across all corpora, meaning that the method is clearly superior when the focus is entirely on grammar.
@@
$\mathcal{C}_{\mathrm{Enron}}$ 0.031 0.039 0.039 -0.017 0.083 4 -4 1 -1
$\mathcal{C}_{\mathrm{Wiki}}$ 0.026 0.003 0.029 0.022 0.036 4 -4 -2 2
$\mathcal{C}_{\mathrm{Stack}}$ 0.005 0.001 -0.002 0.047 -0.044 -5 5 -6 6
$\mathcal{C}_{\mathrm{ACL}}$ 0.057 0.050 0.031 0.106 -0.050 -7 7 -23 23
$\mathcal{C}_{\mathrm{PJ}}$ 0.009 0.005 0.010 0.001 0.019 3 -3 0 0
$\mathcal{C}_{\mathrm{Apric}}$ 0.077 0.042 0.084 0.038 0.129 22 -22 -4 4
$\mathcal{C}_{\mathrm{Trip}}$ 0.037 0.049 0.048 0.022 0.071 17 -17 -1 1
$\mathcal{C}_{\mathrm{Yelp}}$ 0.057 0.046 0.063 0.043 0.083 20 -20 -7 7
$\mathcal{C}_{\mathrm{IMDB}}$ 0.080 0.101 0.091 0.070 0.108 87 -87 -42 42
$\mathcal{C}_{\mathrm{Blog}}$ 0.063 0.038 0.066 0.048 0.083 75 -75 -38 38
$\mathcal{C}_{\mathrm{Amazon}}$ 0.046 0.030 0.046 0.041 0.050 60 -60 -49 49
$\mathcal{C}_{\mathrm{AllNews}}$ 0.043 0.049 0.059 0.034 0.076 101 -101 -15 15
Table 3: Differences in evaluation results when applying POSNoise in LambdaG compared to omitting it. Positive values (shaded in blue) indicate results improvement, while negative values (shaded in red) indicate results degradation by using POSNoise.
$\mathcal{C}_{\mathrm{Enron}}$ 0.010 0.010 0.014 -0.019 0.042 2 -2 1 -1
$\mathcal{C}_{\mathrm{Wiki}}$ 0.026 0.014 0.027 0.036 0.018 2 -2 -4 4
$\mathcal{C}_{\mathrm{Stack}}$ $0$0.004 -0.001 0.008 -0.009 -1 1 -1 1
$\mathcal{C}_{\mathrm{ACL}}$0.022 -0.014 -0.019 -0.033 -0.007 -1 1 5 -5
$\mathcal{C}_{\mathrm{PJ}}$ 0.016 0.000 0.017 0.001 0.032 5 -5 0 0
$\mathcal{C}_{\mathrm{Apric}}$0.003 -0.004 -0.003 $0$0.006 -1 1 0 0
$\mathcal{C}_{\mathrm{Trip}}$ 0.014 0.010 0.019 0.007 0.029 7 -7 0 0
$\mathcal{C}_{\mathrm{Yelp}}$0.014 -0.010 -0.019 -0.003 -0.038 -9 9 -2 2
$\mathcal{C}_{\mathrm{IMDB}}$ 0.015 0.013 0.014 0.019 0.010 8 -8 -16 16
$\mathcal{C}_{\mathrm{Blog}}$ 0.006 0.002 0.006 0.005 0.006 6 -6 -5 5
$\mathcal{C}_{\mathrm{Amazon}}$0.002 -0.001 -0.003 0.003 -0.009 -10 10 -5 5
$\mathcal{C}_{\mathrm{AllNews}}$ 0.027 0.031 0.040 0.015 0.057 76 -76 5 -5
Table 4: Differences in evaluation results when applying POSNoise in LambdaG compared to applying the POSNoise without POS-Tagging (i. e., using a simple star). Positive values (shaded in blue) indicate results improvement, while negative values (shaded in red) indicate results degradation by using full POSNoise.
Secondly, we evaluated LambdaG using a variant of POSNoise similar to TextDistortion, but instead of utilizing POS tags, we replaced all non-functional items with asterisks. The results, presented in Table 4, indicate that the POS tag labels contribute very little information overall, implying that a POS tagger might not be necessary.
As explained in Section 6.3, the advantage of LambdaG is that it can be better interpreted than other methods. The analyst can produce a ranking of all sentences in $\mathcal{D}_{\mathcal{U}}$ and each token can be color-coded according to its value of $\lambda_{G}$, similarly to COAV. Some of these visualizations explaining the features discovered are shown in Section 7.